# MimicMotion: High-Quality Human Motion Video Generation with Confidence-aware Pose Guidance

### v1

@@ Abstract
In recent years, generative artificial intelligence has achieved significant advancements in the field of image generation, spawning a variety of applications. However, video generation still faces considerable challenges in various aspects, such as controllability, video length, and richness of details, which hinder the application and popularization of this technology. In this work, we propose a controllable video generation framework, dubbed MimicMotion, which can generate high-quality videos of arbitrary length mimicking specific motion guidance. Compared with previous methods, our approach has several highlights. Firstly, we introduce confidence-aware pose guidance that ensures high frame quality and temporal smoothness. Secondly, we introduce regional loss amplification based on pose confidence, which significantly reduces image distortion. Lastly, for generating long and smooth videos, we propose a progressive latent fusion strategy. By this means, we can produce videos of arbitrary length with acceptable resource consumption. With extensive experiments and user studies, MimicMotion demonstrates significant improvements over previous approaches in various aspects. Detailed results and comparisons are available on our project page: https://tencent.github.io/MimicMotion.
@@ 1 Introduction
With the rapid development of generative artificial intelligence, video generation is gaining attention in parallel with the growing maturity of image generation techniques. However, video generation is more challenging due to its higher inherent complexities, including the need for high-quality imagery and seamless temporal smoothness. This sets higher standards for video generation technology. In addition to these challenges, controlling the generated content and extension to significant lengths without compromising quality is essential for real-world use. In this paper, we focus on pose-guided video generation conditioned on a reference image. Our goal is to generate a video that not only contains rich imagery details but also adheres to the reference image and the pose guidance.
Currently, there are plenty of works focusing on image-conditioned pose-guided video generation, such as Follow Your Pose [1], DreamPose [2], DisCo [3], MagicDance [4], AnimateAnyone [5], MagicAnimate [6], DreaMoving [7], Champ [8], etc. Though various model architectures and training techniques have been studied for better generation performance, the generated results are unsatisfactory in several aspects. Imagery distortion especially on the regions of human hands is still a common issue which is particularly evident in videos containing large movements. Besides, to achieve good temporal smoothness, imagery details are sometimes sacrificed resulting in videos of blurred frames. In the presence of diverse appearances and motions in videos, accurate pose estimation is inherently challenging. This inaccuracy not only creates a conflict between pose alignment and temporal smoothness but also hinders the model scaling on the training schedule due to overfitting on noisy samples. In addition, due to computational limitations and model capabilities, there are still significant challenges in generating high-quality long videos containing a large number of frames. To solve these problems, we propose a series of approaches for generating long but still smooth videos based on pose guidance and image reference.
To alleviate the negative impact of inaccurate pose estimation, we propose an approach of confidence-aware pose guidance. By introducing the concept of confidence to the pose sequence representation, better temporal smoothness can be achieved and imagery distortion can also be eased. Confidence-based regional loss amplification can make the hand regions more accurate and clear. In addition, we propose a progressive latent fusion method for achieving long but still smooth video generation. Through generating video segments with overlapped frames with the proposed progressive latent fusion, our model can handle arbitrary-length pose sequence guidance. By merging the generated video segments, the final long video can be of good cross-frame smoothness and imagery richness at the same time. For model training, to keep the cost of model training within an acceptable range, our method is based on a generally pre-trained video generation model. The amount of training data is not large and no special manual annotation is required.
@@
    We improve the pose guidance by employing a confidence-aware strategy. In this way, the negative impact of inaccurate pose estimation can be alleviated. This approach not only reduces the influence of noisy samples during training but also corrects erroneous pose guidance during inference.
@@
    While cross-frame overlapped diffusion is a standard technique for generating long videos, we advance it with a position-aware progressive latent fusion approach that improves temporal smoothness at segment boundaries. Extensive experimental results show the effectiveness of the proposed approach.
@@ 2.1 Diffusion models for image/video generation
Diffusion-based models have demonstrated promising results in the fields of image [9, 10, 11, 12, 13] and video generation [14, 15, 16, 17, 18, 19, 20], renowned for their capacity in generative tasks. Diffusion models operating in the pixel domain encounter challenges in generating high-resolution images due to information redundancy and high computational costs. Latent Diffusion Models (LDM) [11] address these issues by performing the diffusion process in low-dimensional latent spaces, significantly enhancing generation efficiency and quality while reducing computational demands. Compared to image generation, video generation demands a more precise understanding of spatial relationships and temporal motion patterns. Recent video generation models leverage diffusion models by adding temporal layers to pre-trained image generation models [14, 15, 21, 22], or utilizing transformers structures [23, 24, 25, 26] to enhance generative capabilities for videos. Stable Video Diffusion (SVD) [20] is one of the most popular open-source models built upon LDM. It offers a straightforward and effective method for image-based video generation and serves as a powerful pre-trained model for this task. Our approach extends SVD for pose-guided video generation, leveraging the pre-trained generative capabilities of SVD.
@@ 2.2 Pose-guided human motion transfer
Pose-to-appearance mapping aims to transfer motion from the source identity to the target identity. Methods based on paired keypoints from source and target images employ local affine transformations [27, 28] or Thin-Plate Spline transformations [29] to warp the source image to match the driving image. These techniques aim to minimize distortion by applying weighted affine transformations, thereby generating poses in the output image that closely resemble those in the driving image. Similarly, methods such as [30, 1, 5, 7] utilize pose stick figures obtained from off-the-shelf human pose detectors as motion indicators and directly generate video frames through generative models. Depth information [7] or 3D human parametric models, such as SMPL (Skinned Multi-Person Linear) [8], can also be used to represent human geometry and motion characteristics from the source video. Nevertheless, these overly dense guidance techniques can rely too much on the signal from the source video, such as the outline of the body, leading to a degradation in the quality of the generated videos, especially when the target identity differs significantly from the source. Our approach, leveraging off-the-shelf human pose detectors, is capable of capturing the motion of the human body in driving videos without introducing excessive extraneous information, thereby ensuring the overall quality of the generated video. Different from existing methods, we introduce confidence-aware pose guidance, which effectively mitigates the influence of inaccurate pose estimation in training and inference. In this way, we achieve superior portrait frame quality, especially in the hand regions.
@@ 2.3 Long video generation
Recent diffusion-based video generation algorithms are constrained to producing videos with durations of only a few seconds, significantly limiting their practical applications. As a result, substantial research efforts have been dedicated to extending the duration of generated videos, leading to the proposal of various approaches to overcome this limitation. Methods like [17, 31] autoregressively predict successive frames, enabling the generation of infinitely long videos. However, these methods often face quality degradation due to error accumulation and the lack of long-term temporal coherence. Hierarchical approach [32, 22] are proposed for generating long videos in a coarse-to-fine manner. It first creates a coarse storyline with keyframes using a global diffusion model, then iteratively refines the video with local diffusion models to produce detailed intermediate frames.
MultiDiffusion [33] combines multiple processes that use pre-trained text-to-image diffusion models to create high-quality images with user-defined controls. It works by applying the model to different parts of an image and using an optimization method to ensure all parts blend seamlessly. This allows users to generate images that meet specific requirements, like certain aspect ratios or spatial layouts, without needing additional training or fine-tuning. Lumiere [34] extends MultiDiffusion to video generation by dividing the video into overlapping temporal segments. Each segment is independently denoised, and an optimization algorithm then combines these denoised segments. This approach ensures high coherence in the generated video, effectively maintaining temporal smoothness across segments. However, our experiments reveal that abrupt transitions can still occur at segment boundaries.
Building upon the principle of MultiDiffusion, we introduce a position-aware progressive latent fusion strategy that enhances temporal smoothness near segment boundaries. We adaptively assign fusion weight based on the temporal position, ensuring a smooth transition at the segment boundaries that further reduces flickering.
@@ 3.1 Preliminaries
A Diffusion Model (DM) learns a diffusion process that generates a probability distribution for a given dataset. In the case of visual content generation tasks, a neural network of DM is trained to reverse the process of adding noise to real data so new data can be progressively generated starting from random noise. For a data sample $\mathbf{x}\sim p_{\text{data}}$ from a specific data distribution $p_{\text{data}}$, the forward diffusion process is defined as a fixed Markov Chain that gradually adds Gaussian noise to the data following:
$q(\mathbf{x_{t}}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$ (1)
for $t=1,\cdots T$, where $T$ is the number of perturbing steps and $\mathbf{x}_{t}$ represents noisy data after adding $t$ steps of noise on the real data $\mathbf{x}_{0}$. This process is controlled by a sequence schedule $\beta_{t}$ which is parameterized by the noising step $t$. Following the closure of normal distribution, $\mathbf{x}_{t}$ can be directly computed with $\mathbf{x}_{0}$ by reforming the above diffusion process as follows:
@@
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{t}=1-\beta_{t}$. Following DDPM [9], a denoising function $\epsilon_{\theta}$ parameterized with $\theta$, commonly implemented with a neural network, is trained by minimizing the mean square error loss as follows:
@@
where $\mathbf{c}$ is an optional condition and $\mathbf{x}_{t}$ is a perturbed version of real data $\mathbf{x}_{0}\sim p_{\text{data}}$ by adding $t$-step noises. In this way, $\epsilon_{\theta}$ can be trained till converge by sampling $\mathbf{x}_{0}$ from real data distribution and a time step $t$, with an optional condition $\mathbf{c}$.
@@ 3.2 Data preparation
To train a pose-guided video diffusion model, we collect a video dataset containing various human motions. Leveraging the powerful capability of the generally pre-trained image-to-video model, the dataset need not be excessively large, as the pre-trained model already has a good prior.
Given a video from our dataset, the training sample is constructed with three parts: a reference image (denoted as $I_{\text{ref}}$), a sequence of raw video frames, and the corresponding poses. Firstly, basic pre-processing operations like frame resizing and cropping are applied to the raw video to get a sequence of video frames with a fixed aspect ratio. For a given video, a fixed number of frames are randomly sampled at equal intervals as input video frames to the diffusion model. The input reference image is randomly sampled from the same video at a location not limited to the sampled video frame. This reference image is pre-processed in the same way as the video frames. Another input of the model is the pose sequence, which is extracted from the video frames with DWPose [35] frame by frame.
@@ 3.3 Pose-guided Video Diffusion Model
The goal of MimicMotion is to generate high-quality, pose-guided human videos from a single reference image and a sequence of poses to mimic. This task involves the synthesis of realistic motion that adheres to the provided pose sequence while maintaining visual fidelity to the reference image. We exploit the ability of a specific pre-trained video diffusion model to reduce the data requirement and computational cost of training a video diffusion model from scratch. Stable Video Diffusion (SVD) [20] is an open-source image-to-video diffusion model trained on a large-scale video dataset. It shows good performance on both video quality and diversity compared with the other contemporary models. The model structure of MimicMotion is designed to integrate a pre-trained Stable Video Diffusion (SVD) model to leverage its image-to-video generation capabilities.
Learning a diffusion process in pixel space is costly and this is more severe in generating high-definition videos involving many frames. We follow the Latent Diffusion Model (LDM) [36] to encode pixels into latent space so diffusion can be conducted in a low-dimension latent space. LDM adopts a pair of autoencoders, consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. Given a data sample $\mathbf{x}$, it is encoded into the latent space as $\mathbf{z}=\mathcal{E}(\mathbf{x})$. Conversely, the latent vector $\mathbf{z}$ can be decoded back into pixel space via $\mathbf{x}=\mathcal{D}(\mathbf{z})$.
@@
Figure 2: MimicMotion integrates an image-to-video diffusion model with novel confidence-aware pose guidance. The model’s trainable components consist of a spatiotemporal U-Net and a PoseNet for introducing pose sequence as the condition. Key features of confidence-aware pose guidance include: 1) The pose sequence condition is accompanied by keypoint confidence scores, enabling the model to adaptively adjust the influence of pose guidance based on the score. 2) The regions with high confidence are given greater weight in the loss function, amplifying their impact in training.
@@
Figure 2 shows the structure of our model. The core structure of our model is a latent video diffusion model with a U-Net for progressive denoising in latent space. The VAE encoder on input video frames and the corresponding decoder for getting denoised video frames are both adopted from SVD and these parameters are frozen. The VAE encoder is applied independently to each frame of the input video as well as to the conditional reference image, operating on a per-frame basis without considering temporal or cross-frame interactions. Differently, the VAE decoder processes the latent features, which undergo spatiotemporal interaction from U-Net. To ensure the generation of a smooth video, the VAE decoder incorporates temporal layers alongside the spatial layers, mirroring the architecture of the VAE encoder.
@@
In addition to the input video frames, the reference image and the sequence of poses are two other inputs of the model. The reference image is fed into the diffusion model along two separate pathways. One pathway involves feeding the image into each block of the U-Net. Specifically, through a visual encoder like CLIP [37], the image feature is extracted and fed into the cross-attention of every U-Net block for finally controlling the output results. The other pathway targets the input latent features. Similar to the raw video frames, the input reference image is encoded with the same frozen VAE encoder to get its representation in the latent space. The latent feature of the single reference image is then duplicated along the temporal dimension to align with the features of input video frames. The duplicated latent reference images are concatenated with latent video frames along the channel dimension and then fed into U-Net for diffusion altogether.
@@
For introducing the guidance of poses, PoseNet, which is implemented with multiple convolution layers, is designed as a trainable module for extracting features of the input sequence of poses. The reason for not using the VAE encoder is that the pixel value distribution of the pose sequence is different from that of common images on which the VAE autoencoder is trained. With PoseNet, the features of poses are extracted and then element-wisely added to the output of the first convolution layer of U-Net. In this way, the influence of the posture guidance can take effect from the very beginning of denoising. We do not add pose guidance to every U-Net block for the following considerations: a) the sequence pose is extracted frame by frame without any temporal interaction so it may confuse the spatio-temporal layers within U-Net when it takes effect on these layers directly; b) excessive involvement of the pose sequence may degrade the performance of the pre-trained image-to-video model.

### v2

@@ Abstract
In recent years, while generative AI has advanced significantly in image generation, video generation continues to face challenges in controllability, length, and detail quality, which hinder its application. We present MimicMotion, a framework for generating high-quality human videos of arbitrary length using motion guidance. Our approach has several highlights. Firstly, we introduce confidence-aware pose guidance that ensures high frame quality and temporal smoothness. Secondly, we introduce regional loss amplification based on pose confidence, which reduces image distortion in key regions. Lastly, we propose a progressive latent fusion strategy to generate long and smooth videos. Experiments demonstrate the effectiveness of our approach in producing high-quality human motion videos. Videos and comparisons are available at https://tencent.github.io/MimicMotion.
Machine Learning, ICML
@@ 1 Introduction
Video generation has advanced alongside image generation. However, video generation is more challenging due to its higher inherent complexities, including the need for high-quality imagery and seamless temporal smoothness. In addition, controlling the generated content and extension to significant lengths is essential for real-world use. In this paper, we focus on pose-guided video generation with a reference image. Our goal is to generate a video that contains rich imagery details and adheres to the reference image and the pose guidance.
Currently, there are plenty of works focusing on this task, such as DisCo (Wang et al., 2023), AnimateAnyone (Hu, 2024), MagicDance (Chang et al., 2023), MagicAnimate (Xu et al., 2024), etc. Though various techniques have been studied, the results are unsatisfactory in several aspects. Imagery distortion especially on the regions of human hands is still a common issue which is particularly evident in videos containing large movements. Besides, to achieve good temporal smoothness, imagery details are sometimes sacrificed resulting in videos of blurred frames. In the presence of diverse appearances and motions in videos, accurate pose estimation is inherently challenging. This inaccuracy not only creates a conflict between pose alignment and temporal smoothness but also hinders the model scaling on the training schedule due to overfitting on noisy samples. In addition, due to computational limitations and model capabilities, there are still significant challenges in generating high-quality long videos containing a large number of frames. To solve these problems, we propose a series of approaches for generating long but still smooth videos based on pose guidance and image reference.
To alleviate the negative impact of inaccurate pose estimation, we propose an approach of confidence-aware pose guidance. By introducing the concept of confidence to the pose sequence representation, better temporal smoothness can be achieved and imagery distortion can also be eased. Confidence-based regional loss amplification can make the hand regions more accurate and clear. In addition, we propose a progressive latent fusion method for achieving long but still smooth video generation. Through generating video segments with overlapped frames with the proposed progressive latent fusion, our model can handle arbitrary-length pose sequence guidance. By merging the generated video segments, the final long video can be of good cross-frame smoothness and imagery richness at the same time (see Figure 1). For model training, to keep the cost of model training within an acceptable range, our method is based on a generally pre-trained video generation model. The amount of training data is not large and no special manual annotation is required.
@@
    We improve the pose guidance by employing a confidence-aware strategy. In this way, we alleviate the negative impact of inaccurate pose estimation. This approach not only reduces the influence of noisy samples during training but also corrects erroneous pose guidance during inference.
@@
    While overlapped diffusion is a standard technique for generating long videos, we advance it with position-aware progressive latent fusion that improves temporal smoothness at segment boundaries. Extensive experiments show the effectiveness of the proposed approach.
@@ 2.1 Diffusion Models for Image/Video Generation
Diffusion-based models have demonstrated promising results in the fields of image (Ho et al., 2020; Song et al., 2020; Rombach et al., 2022a; Zhang et al., 2023; Wang et al., 2024a) and video generation (Singer et al., 2022; Guo et al., 2023; Ho et al., 2022; He et al., 2022; Chen et al., 2023, 2024; Blattmann et al., 2023), renowned for their capacity in generative tasks. Diffusion models operating in the pixel domain encounter challenges in generating high-resolution images due to information redundancy and high computational costs. Latent Diffusion Models (LDM) (Rombach et al., 2022a) address these issues by performing the diffusion process in low-dimensional latent spaces, significantly enhancing generation efficiency and quality while reducing computational demands. Recently, efforts have been made to improve training efficiency through noise schedule (Zheng et al., 2024a, b). Compared to image generation, video generation demands a more precise understanding of spatial relationships and temporal motion patterns. Recent video generation models leverage diffusion models by adding temporal layers to pre-trained image generation models (Singer et al., 2022; Guo et al., 2023; Wu et al., 2023; Wang et al., 2024c), or utilizing DiT (Peebles & Xie, 2023) structures to enhance generative capabilities for videos (Yan et al., 2021; Yu et al., 2023; Ma et al., 2024a; Bao et al., 2024). Stable Video Diffusion (SVD) (Blattmann et al., 2023) is one of the most popular open-source models built upon LDM. It offers a straightforward and effective method for image-based video generation and serves as a powerful pre-trained model for this task. Our approach extends SVD for pose-guided video generation, leveraging its pre-trained generative capabilities.
@@ 2.2 Pose-Guided Human Motion Transfer
Pose-to-appearance mapping aims to transfer motion from the source identity to the target identity. Methods based on paired keypoints from source and target images employ local affine transformations (Siarohin et al., 2019, 2021) or Thin-Plate Spline transformations (Zhao & Zhang, 2022) to warp the source image to match the driving image. These techniques aim to minimize distortion by applying weighted affine transformations, thereby generating poses in the output image that closely resemble those in the driving image. Similarly, methods such as (Chan et al., 2019; Tu et al., 2024; Ma et al., 2024b; Hu, 2024; Feng et al., 2023; Peng et al., 2024; Li et al., 2024) utilize pose stick figures obtained from off-the-shelf human pose detectors as motion indicators and directly generate video frames through generative models. Depth information (Feng et al., 2023) or 3D human parametric models, such as SMPL (Skinned Multi-Person Linear) (Zhu et al., 2024; Wang et al., 2024b), can also be used to represent human geometry and motion characteristics from the source video. Nevertheless, these overly dense guidance techniques can rely too much on the signal from the source video, such as the outline of the body, leading to a degradation in the quality of the generated videos, especially when the target identity differs significantly from the source. Our approach, leveraging off-the-shelf human pose detectors, is capable of capturing the motion of the human body in driving videos without introducing excessive extraneous information, thereby ensuring the overall quality of the generated video. In addition, we introduce confidence-aware pose guidance, which effectively mitigates the influence of inaccurate pose estimation in training and inference. In this way, we achieve superior portrait frame quality, especially in the hand regions.
@@ 2.3 Long video generation
Recent diffusion-based video generation algorithms are constrained to producing videos with durations of only a few seconds, significantly limiting their practical applications. As a result, substantial research efforts have been dedicated to extending the duration of generated videos, leading to the proposal of various approaches to overcome this limitation. Methods like (He et al., 2022; Voleti et al., 2022) autoregressively predict successive frames, enabling the generation of infinitely long videos. However, these methods often face quality degradation due to error accumulation and the lack of long-term temporal coherence. Hierarchical approach (Yin et al., 2023; Wang et al., 2024c) are proposed for generating long videos in a coarse-to-fine manner. It first creates a coarse storyline with keyframes using a global diffusion model, then iteratively refines the video with local diffusion models to produce intermediate frames.
MultiDiffusion (Bar-Tal et al., 2023) combines multiple processes that use pre-trained text-to-image diffusion models to create high-quality images with user-defined controls. It works by applying the model to different parts of an image and using an optimization method to ensure all parts blend seamlessly. This allows users to generate images that meet specific requirements, like certain aspect ratios or spatial layouts, without needing additional training or fine-tuning. Lumiere (Bar-Tal et al., 2024) extends MultiDiffusion to video generation by dividing the video into overlapping temporal segments. Each segment is independently denoised, and an optimization algorithm then combines these denoised segments. This approach ensures high coherence in the generated video, effectively maintaining temporal smoothness across segments. However, our experiments reveal that abrupt transitions can still occur at segment boundaries.
Building upon this principle, we introduce a position-aware progressive latent fusion strategy that enhances temporal smoothness near segment boundaries. We adaptively assign fusion weight based on the temporal position, ensuring a smooth transition at the segment boundaries.
@@ 3.1 Preliminaries
A Diffusion Model (DM) learns a diffusion process that generates a probability distribution for a given dataset. In the case of visual content generation tasks, a neural network of DM is trained to reverse the process of adding noise to real data so new data can be progressively generated starting from random noise. For a data sample $\mathbf{x}\sim p_{\text{data}}$ from a specific data distribution $p_{\text{data}}$, the forward diffusion process is defined as a fixed Markov Chain that gradually adds Gaussian noise to the data, following: $q(\mathbf{x_{t}}\mid\mathbf{x}_{t-1})=\mathcal{N}(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I})$ for $t=1,\cdots T$, where $T$ is the number of perturbing steps and $\mathbf{x}_{t}$ represents noisy data after adding $t$ steps of noise on the real data $\mathbf{x}_{0}$. This process is controlled by a schedule $\beta_{t}$. The above diffusion process can be reformed as:
@@
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ and $\alpha_{t}=1-\beta_{t}$. Following DDPM (Ho et al., 2020), a denoising function $\epsilon_{\theta}$ with parameter $\theta$ is trained by minimizing the mean squared error:
@@
where $\mathbf{c}$ is the condition. We then train $\epsilon_{\theta}$ until convergence.
@@ 3.2 Data Preparation
To train a pose-guided video diffusion model, we collect a video dataset containing various human motions. The dataset need not be excessively large, as the pre-trained model already has a good prior.
Given a video from our dataset, the training sample is constructed with three parts: a reference image (denoted as $I_{\text{ref}}$), a sequence of raw video frames, and the corresponding poses. Basic pre-processing operations of resizing and cropping are applied to the raw video to get a sequence of video frames. For a given video, a fixed number of frames are randomly sampled at equal intervals as input video frames to the diffusion model. The input reference image is independently sampled from the video and pre-processed in the same way as the video frames. Another input of the model is the pose sequence, which is extracted from the video frames with DWPose (Yang et al., 2023) frame by frame.
@@ 3.3 Pose-Guided Video Diffusion Model
Figure 2: MimicMotion integrates an image-to-video diffusion model with confidence-aware pose guidance, utilizing a spatiotemporal U-Net and a PoseNet for introducing pose sequence as the condition. Our design features adaptive adjustment of pose influence via keypoint confidence scores and prioritizes high-confidence hand regions in the loss function.
The goal of MimicMotion is to generate high-quality, pose-guided human videos from a single reference image and a sequence of poses to mimic. This task involves the synthesis of realistic motion that adheres to the provided pose sequence while maintaining visual fidelity to the reference image. We exploit the ability of a specific pre-trained video diffusion model to reduce the data requirement and computational cost of training a video diffusion model from scratch. Stable Video Diffusion (SVD) (Blattmann et al., 2023) is an open-source image-to-video diffusion model trained on a large-scale video dataset. It shows good performance on both video quality and diversity compared with the other contemporary models. The model structure of MimicMotion is designed to integrate a pre-trained Stable Video Diffusion (SVD) model to leverage its image-to-video generation capabilities.
Learning a diffusion process in pixel space is costly, especially for high-definition videos involving many frames. We follow the Latent Diffusion Model (LDM) (Rombach et al., 2022b) to encode pixels into a lower-dimensional latent space. LDM adopts a pair of autoencoders, consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. Given a data sample $\mathbf{x}$, it is encoded into the latent space as $\mathbf{z}=\mathcal{E}(\mathbf{x})$. Conversely, the latent vector $\mathbf{z}$ can be decoded into pixel space via $\mathbf{x}=\mathcal{D}(\mathbf{z})$.
@@
Figure 2 shows the structure of our model. The VAE encoder on input video frames and the corresponding decoder for getting denoised video frames are both adopted from SVD and these parameters are frozen. The VAE encoder is applied independently to each frame of the input video and the conditional reference image. In contrast, the VAE decoder processes the latent features that undergo spatiotemporal interaction from U-Net. To enhance the temporal smoothness of low-level details, the VAE decoder incorporates temporal layers alongside the spatial layers.
@@
The reference image and the sequence of poses are two other inputs of the model. The reference image follows two pathways: First, it is processed through CLIP (Radford et al., 2021) to extract features that are fed into each U-Net block’s cross-attention. Second, the image is encoded by the frozen VAE encoder to obtain its latent representation, which is then temporally duplicated to match the video frame dimensions. This latent representation is concatenated with the video frame features along the channel dimension before entering the U-Net.
@@
For pose guidance, PoseNet consisting of multiple convolution layers is designed to extract features from input pose sequences. The reason for not using the VAE encoder is that the pixel value distribution of the pose sequence is different from that of common images on which the VAE autoencoder is trained. The features of poses are then element-wisely added to the output of the first convolution layer of U-Net. Pose guidance is not added to every U-Net block due to: a) the sequence pose being extracted without temporal interaction may confuse the spatio-temporal layers within U-Net; b) excessive involvement of the pose sequence may degrade the performance of the pre-trained image-to-video model.